A person drops a stone from a top of the bridge that is 15m high. How long in seconds will it take the stone to reach the water?


1
Expert's answer
2021-03-24T19:39:31-0400

"h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot 15/9.81}=1.75\ (s)"